The following table relates to the tourist arrivals during 1990 to 1996 in Botswana:

Years:                    1990   1991    1992   1993  1994  1995  1996

Tourist’s arrivals:    18       20        23       25      24      28       30

Fit a straight line trend by the method of least squares and estimates the number of tourists that would arrives in the year 2000.



1
Expert's answer
2021-12-05T18:25:41-0500

Let "X=" the number of years since 1990


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 y & 18 & 20 & 23 & 25 & 24 & 28 & 30 \\
 
\end{array}"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & x & y & xy & x^2 &y^2 \\ \hline
 & 0 & 18 & 0 & 0 & 324 \\
 \hdashline
 & 1 & 20 & 20 & 1 & 400 \\
 \hdashline
 & 2 & 23 & 46 & 4 & 529 \\
 \hdashline
 & 3 & 25 & 75 & 9 & 625 \\
 \hdashline
 & 4 & 24 & 96 & 16 & 576 \\
 \hdashline
 & 5 & 28 & 140 & 25 & 784 \\
 \hdashline
 & 6 & 30 & 180 & 36 & 900 \\
 \hdashline
 Sum= & 21 & 168 & 557 & 91 & 4138 \\
 \hdashline
\end{array}"

"\bar{x}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nx_i=\dfrac{21}{7}=3"

"\bar{y}=\dfrac{1}{n}\displaystyle\sum_{i=1}^ny_i=\dfrac{168}{7}=24"

"SS_{xx}=\displaystyle\sum_{i=1}^nx_i^2-\dfrac{1}{n}\bigg(\displaystyle\sum_{i=1}^nx_i\bigg)^2=91-\dfrac{21^2}{7}=28"

"SS_{yy}=\displaystyle\sum_{i=1}^ny_i^2-\dfrac{1}{n}\bigg(\displaystyle\sum_{i=1}^ny_i\bigg)^2=4138-\dfrac{168^2}{7}=106"

"SS_{xy}=\displaystyle\sum_{i=1}^nx_i\displaystyle\sum_{i=1}^ny_i-\dfrac{1}{n}\bigg(\displaystyle\sum_{i=1}^nx_i\bigg)\bigg(\displaystyle\sum_{i=1}^nx_i\bigg)"

"=557-\dfrac{21(168)}{7}=53"

Based on the above calculations, the slope "m" and the y-intercept "n" are obtained as follows:


"m=\dfrac{SS_{xy}}{SS_{xx}}=\dfrac{53}{28}=1.8929"

"y=\bar{y}-m\bar{x}=24-\dfrac{53}{28}(3)=\dfrac{513}{28}=18.3214"

So the trend line is


"y=18.3214+1.8929x"

The number of tourists that would arrives in the year 2000 is


"y=18.3214+1.8929(10)=37"